Circular No. 6685

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMETS C/1997 L3 AND C/1997 L4 C. St. Cyr, Naval Research Laboratory, reports, for the SOHO-LASCO

Consortium (which consists of the Naval Research Laboratory, Laboratoire d’Astronomie Spatiale, Max-Planck-Institut für Aeronomie, and the University of Birmingham), the discovery of two more comets in images obtained by the LASCO white-light coronagraphs C2 and C3 aboard the ESA/NASA Solar and Heliospheric Observatory (SOHO) spacecraft. Both comets are of mag 4–5 and show faint tails. The first comet, C/1997 L3, was discovered by Shane Stezelberger, and the second, C/1997 L4, was discovered by Darren Lewis. The following approximate discovery positions are summarized from the more complete reductions by G. V. Williams on MPEC 1997-M07 and 1997-M08 from measurements of apparent solar distance and position angle by St. Cyr and Lewis:

1997 UT  α_{2000}  δ_{2000}
June 12.523  5^{h}18.^{m}7  +19^{o}01′  C/1997 L3
June 14.355  5 28.9  +19 02  C/1997 L4

The comets were observed until June 13.631 and 15.771, respectively. St. Cyr suggests that both comets are Kreutz sungrazers, although he notes that there is an 8^{o} difference in the position angles of the comets at comparable elongations from the center of the sun. Orbital computations by the undersigned confirm that both comets are indeed Kreutz sungrazers, but that the two objects, although at perihelion only two days apart (on June 13.85 and 15.91, respectively), seem to belong to different subgroups (Marsden 1967, A.J. 72, 1170; 1989, ibid. 98, 2306), immediately suggesting that the number of small Kreutz members must be very large indeed and the process of producing new members from those that survive perihelion passage almost continuous. It appears that C/1997 L3 belongs to Subgroup II (which includes C/1882 R1 and C/1965 S1), while C/1997 L4 belongs to Subgroup I (which includes C/1843 D1 and C/1963 R1). Further details are on the MPEC s cited above.

COMET C/1995 O1 (HALE-BOPP) Total visual magnitude estimates: May 26.35 UT, 1.3 (T. Lovejoy,

Brisbane, Queensland, Australia; naked eye); 29.90, 1.6 (R. Lourençon, Jundiai, Brazil, naked eye); June 2.43, 2.0 (A. Pearce, Subiaco, Western Australia, naked eye); 7.88, 2.1 (J. G. de S. Aguiar, Campinas, Brazil, 11 × 80 binoculars); 14.44, 2.7 (Pearce, City Beach, Western Australia; 20 × 80 binoculars); 17.44, 2.8 (Pearce).

1997 June 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden